OpenStudy (anonymous):

The Haul it all truck rental charges $65 for a two day rental, plus $0.15 for each mile over 100 miles. Burt's total charge for a two day rental was $83.75. How many miles did he drive?

OpenStudy (darkbluechocobo):

Alright yay another one :P

OpenStudy (darkbluechocobo):

Basically you know the total price is 83.75

OpenStudy (darkbluechocobo):

so automatically you can subtract 83.75 by 65 which is 18.75

OpenStudy (darkbluechocobo):

So now you know how much he payed over for the miles extended over 100

OpenStudy (anonymous):

I got 125 miles for an answer :) is that right?

OpenStudy (darkbluechocobo):

xD one moment still trying to figure out the end

OpenStudy (darkbluechocobo):

Yes you got it :3

OpenStudy (darkbluechocobo):

Nice job :p

OpenStudy (anonymous):

Thank you soooo much!!! Your such a big help! :D

OpenStudy (darkbluechocobo):

it has been a pleasure helping you out again :p If you ever need help just give me a message xD

OpenStudy (anonymous):

Thank you!!!!

OpenStudy (darkbluechocobo):

You are very welcome my friend

OpenStudy (anonymous):

I have another problem ;) The width of a rectangle is 1 foot shorter than half the length. The perimeter of the rectangle is 64 feet. What are the dimensions of the rectangle?

OpenStudy (darkbluechocobo):

o god

OpenStudy (darkbluechocobo):

geometry

OpenStudy (anonymous):

:)

OpenStudy (darkbluechocobo):

so basically the total is 64 because its the perimeter of the whole rectangle

OpenStudy (darkbluechocobo):

alright then ┻━┻ ︵ヽ(`Д´)ノ︵ ┻━┻

OpenStudy (darkbluechocobo):

so its blankity blank=64 at the current moment

OpenStudy (anonymous):

Hahaha! That's exactly what I have!!! :D

OpenStudy (darkbluechocobo):

XD

OpenStudy (darkbluechocobo):

you are going to have to use l and w

OpenStudy (darkbluechocobo):

so w is one foot shorter than half the length

OpenStudy (darkbluechocobo):

w-.5l=64 maybe

OpenStudy (darkbluechocobo):

.5l-w=64

OpenStudy (anonymous):

How would you solve for l and w? You need to get both of them alone, right?

OpenStudy (darkbluechocobo):

well it gives you that w is 1 foot shorter than the length so not really

OpenStudy (darkbluechocobo):

you have to work them together

OpenStudy (anonymous):

Oh! Okay :)

OpenStudy (darkbluechocobo):

Wait

OpenStudy (darkbluechocobo):

I am an idiot ._. I didn't work in the 1 foot

OpenStudy (darkbluechocobo):

.5l-1=64

OpenStudy (anonymous):

Oh that makes a lot more sense! :) thank you!

OpenStudy (darkbluechocobo):

basically now that you know that do you know how to find the perimeter of a rectangle?

OpenStudy (darkbluechocobo):

Basically 64= 2l+2w

OpenStudy (darkbluechocobo):

so since its 1 foot shorter than half the length you substitute that equation with W

OpenStudy (darkbluechocobo):

so 64=2l+2(.5l-1)

OpenStudy (darkbluechocobo):

wait i got mixed up

OpenStudy (darkbluechocobo):

so 64=2l+1l-2

OpenStudy (darkbluechocobo):

64+2 is 66 so 66=3l

OpenStudy (darkbluechocobo):

L= 22

OpenStudy (darkbluechocobo):

To find W you plug it into that equation

OpenStudy (anonymous):

So then you would subtract one from 11?

OpenStudy (darkbluechocobo):

W=.5l -1

OpenStudy (darkbluechocobo):

? where did you get 11 lol

OpenStudy (anonymous):

22+22+10+10=64

OpenStudy (darkbluechocobo):

XD boom you got it L=22 W=10 *claps*

OpenStudy (anonymous):

YAY!! Thank you so much your seriously a life saver!!!!!

OpenStudy (darkbluechocobo):

xD You are welcome

OpenStudy (darkbluechocobo):

:p Anytime you can get a hold of me for a question I will try to do the best to answer it